OpenStudy (russianboss1):

which expression is equal to (^5sqrt x^6+9)^4 algebra 2 please help someone don't understand

OpenStudy (freckles):

\[(\sqrt[5]{x^6}+9)^4 ?\]

OpenStudy (freckles):

if so you can expand it: \[(a+b)^4=a^4+4a^3b+6a^2b^2+4ab^3+b^4\]

OpenStudy (russianboss1):

never mind already got the answer and explanation already thanks

OpenStudy (freckles):

k